# Horizon Brightened Acceleration Radiation from Massive Vector Fields

### v1

@@ IV Detector models and atom-field interaction
We model the atom as a two-level system with gap $\omega>0$ moving on the infalling geodesic specified in Section II. The Proca field is expanded in properly normalized outgoing Schwarzschild modes as in Section III. Interaction is treated perturbatively in the coupling constants. We consider two couplings: a charged-monopole (current) coupling, where the atomic monopole couples to $u^{\mu}A_{\mu}$ along the trajectory, and a physical dipole coupling to the local electric field. In each case, excitation of a ground-state atom in the Boulware vacuum is mediated by the counter-rotating channel (atom excitation + field emission). The resulting amplitudes reduce to a single worldline integral with the near-horizon phase $\nu[t(\tau)-r_{*}(\tau)]$.
@@ IV.1 Charged-monopole (current) coupling
$L_{\text{int}}(\tau)=g_{V}\,m(\tau)\,u^{\mu}(\tau)A_{\mu}\!\left(x(\tau)\right),$ (35)
@@
$m(\tau)=m_{eg}\,e^{+i\omega\tau}\,\sigma_{eg}+m_{ge}\,e^{-i\omega\tau}\,\sigma_{ge},\qquad m_{ge}=m_{eg}^{\ast},$ (36)
@@
$A_{\mu}(x)=\sum_{\ell m}\sum_{\lambda=1}^{3}\int_{m_{V}}^{\infty}\!d\nu\;\left[a_{\nu\ell m\lambda}\,\mathcal{A}^{(\lambda)}_{\mu;\nu\ell m}(x)+a_{\nu\ell m\lambda}^{\dagger}\,\mathcal{A}^{(\lambda)\,*}_{\mu;\nu\ell m}(x)\right],$ (37)
@@
$\mathcal{A}^{(\lambda)}_{\mu;\nu\ell m}(x)\sim\frac{1}{r}\,\frac{\varepsilon^{(\lambda)}_{\mu}(\nu\,\hat{\mathbf{k}})}{\sqrt{k_{\infty}(\nu)}}\,Y_{\ell m}(\theta,\phi)\,e^{-i\nu t}\,e^{-ik_{\infty}(\nu)r_{*}},\qquad k_{\infty}(\nu)=\sqrt{\nu^{2}-m_{V}^{2}}.$ (38)
Near the horizon each outgoing mode reduces to the universal WKB form $e^{-i\nu[t-r_{*}]}$, consistent with (15) and the sign convention fixed earlier.
Starting from the ground state $|g;0_{B}\rangle$, the only first-order process that excites the atom in the Boulware vacuum is the counter-rotating one, $|g;0_{B}\rangle\to|e;1_{\nu\ell m\lambda}\rangle$. The corresponding amplitude for a selected cavity mode $(\nu,\ell,m,\lambda)$ is
$\mathcal{A}^{\text{(mono)}}_{\nu\ell m\lambda}=-\,ig_{V}m_{eg}\!\int_{-\infty}^{+\infty}\!d\tau\;e^{+i\omega\tau}\,u^{\mu}(\tau)\,\mathcal{A}^{(\lambda)\,*}_{\mu;\nu\ell m}\!\left(x(\tau)\right).$ (39)
Inserting (38) and restricting to the outgoing branch selected by the cavity, we obtain a one-dimensional integral of the form
$\mathcal{A}^{\text{(mono)}}_{\nu\ell m\lambda}=-\,ig_{V}m_{eg}\!\int d\tau\;\frac{u^{\mu}(\tau)\,\varepsilon^{(\lambda)}_{\mu}(\nu\,\hat{\mathbf{k}})}{\sqrt{k_{\infty}(\nu)}\,r(\tau)}\,Y_{\ell m}^{\ast}\!\left(\theta(\tau),\phi(\tau)\right)\,e^{\,i\Phi(\tau,\nu)},$ (40)
@@
$\Phi(\tau,\nu)\equiv\omega\tau+\nu\,t(\tau)+k_{\infty}(\nu)\,r_{*}(\tau).$ (41)
Using the geodesic relations (5)-(8) and the near-horizon WKB replacement $k_{\infty}(\nu)\to\nu$ inside the rapidly varying part, we may equivalently write
$\Phi(\tau,\nu)=\omega\tau+\nu\,\left[\,t(\tau)-r_{*}(\tau)\,\right]+\Delta(\nu)\,r_{*}(\tau),\qquad\Delta(\nu)\equiv\nu-k_{\infty}(\nu),$ (42)
where $\Delta(\nu)$ is smooth and slowly varying. The polarization contraction can be organized with the projector defined in (33). Summing over the three Proca polarizations gives the smooth weight
$\sum_{\lambda=1}^{3}\big|u^{\mu}\varepsilon^{(\lambda)}_{\mu}\big|^{2}=u^{\mu}u^{\nu}\left(-\eta_{\mu\nu}+\frac{k_{\mu}k_{\nu}}{m_{V}^{2}}\right)=-\,u^{\mu}u_{\mu}+\frac{(u\!\cdot\!k)^{2}}{m_{V}^{2}}=-1+\frac{\Omega^{2}(\tau,\nu)}{m_{V}^{2}},$ (43)
where $\Omega(\tau,\nu)=-k_{\mu}u^{\mu}$ is the comoving frequency along the worldline, equal to the redshift-Doppler factor given in (15) for the near-horizon outgoing branch.
@@
$P_{\text{exc}}^{\text{(mono)}}(\nu,\ell)=\sum_{m,\lambda}\Big|\mathcal{A}^{\text{(mono)}}_{\nu\ell m\lambda}\Big|^{2}=g_{V}^{2}\,|m_{eg}|^{2}\,\Bigg|\int d\tau\,\frac{e^{\,i\Phi(\tau,\nu)}}{\sqrt{k_{\infty}(\nu)}\,r(\tau)}\,\mathcal{S}_{\ell m}(\tau,\nu)\Bigg|^{2},$ (44)
@@
$\mathcal{S}_{\ell m}(\tau,\nu)\equiv Y_{\ell m}^{\ast}\!\left(\theta(\tau),\phi(\tau)\right)\,\left[\,-1+\frac{\Omega^{2}(\tau,\nu)}{m_{V}^{2}}\,\right]^{1/2}$ (45)
collects the angular dependence and the polarization weight. In the large-$\omega$ regime and with the cavity isolating a single outgoing channel, the integral in (44) will be evaluated by stationary phase around the near-horizon segment where $\Phi^{\prime}(\tau,\nu)=0$, with the rapidly varying part governed by $\nu[t(\tau)-r_{*}(\tau)]$ and the smooth prefactor supplied by $\mathcal{S}_{\ell m}$ and $1/r(\tau)$. The appearance of $\Omega(\tau,\nu)$ in (45) traces directly to the current coupling $u\!\cdot\!a$ and encodes the distinctive mass/polarization dependence of the vector response.
@@ IV.2 Physical dipole coupling
$L_{\text{int}}(\tau)\;=\;-\,d^{\mu}(\tau)\,F_{\mu\nu}\!\left(x(\tau)\right)\,u^{\nu}(\tau),\qquad u_{\mu}d^{\mu}(\tau)=0,$ (46)
@@
$d^{\mu}(\tau)=d_{eg}\,e^{+i\omega\tau}\,\hat{d}^{\mu}\,\sigma_{eg}+d_{ge}\,e^{-i\omega\tau}\,\hat{d}^{\mu}\,\sigma_{ge},\qquad\hat{d}^{\mu}\hat{d}_{\mu}=-1,\quad u_{\mu}\hat{d}^{\mu}=0.$ (47)
Using the properly normalized outgoing modes from Section III, the field strength associated with a single mode $(\nu,\ell,m,\lambda)$ is
$F_{\mu\nu}^{(\lambda)}(x)\;=\;i\left[k_{\mu}\,\mathcal{A}^{(\lambda)}_{\nu;\nu\ell m}(x)-k_{\nu}\,\mathcal{A}^{(\lambda)}_{\mu;\nu\ell m}(x)\right],$ (48)
where $\mathcal{A}^{(\lambda)}_{\mu;\nu\ell m}$ has the asymptotics in (38) and $k^{\mu}=(\nu,\mathbf{k})$ with $k_{\infty}=\sqrt{\nu^{2}-m_{V}^{2}}$. Contracting with $u^{\nu}$ defines the electric field in the atom’s instantaneous rest frame scully1997quantum; cohen1998atom,
$E_{\mu}\;\equiv\;F_{\mu\nu}\,u^{\nu}\;=\;i\left[-(u\!\cdot\!k)\,\mathcal{A}_{\mu}+(u\!\cdot\!\mathcal{A})\,k_{\mu}\right].$ (49)
@@
$E_{\mu}=i\frac{Y_{\ell m}}{r\sqrt{k_{\infty}(\nu)}}\Big[k_{\mu}(\varepsilon\cdot u)-\Omega(\tau,\nu)\varepsilon_{\mu}\Big]e^{-i\nu t}e^{-ik_{\infty}r_{*}}+\cdots,\qquad\Omega=-k\cdot u.$ (50)
The counter-rotating channel $|g;0_{B}\rangle\!\to\!|e;1_{\nu\ell m\lambda}\rangle$ gives the first-order amplitude
$\mathcal{A}^{\text{(dip)}}_{\nu\ell m\lambda}=-\,ig_{d}d_{eg}\!\int d\tau\;\hat{d}^{\mu}\,E_{\mu}^{(\lambda)\,*}\!\left(x(\tau)\right)\,e^{+i\omega\tau},$ (51)
which, after inserting (50) and using the cavity-selected outgoing branch, reduces to
$\mathcal{A}^{\text{(dip)}}_{\nu\ell m\lambda}=-g_{d}d_{eg}\int d\tau\frac{e^{i\Phi(\tau,\nu)}}{\sqrt{k_{\infty}(\nu)}r(\tau)}\left[\Omega(\tau,\nu)\hat{d}\cdot\varepsilon^{(\lambda)}-(\hat{d}\cdot k)(\varepsilon^{(\lambda)}\cdot u)\right]Y_{\ell m}^{\ast}(\theta(\tau),\phi(\tau)),$ (52)
with the same phase used in (41)-(42),
$\Phi(\tau,\nu)=\omega\tau+\nu\,t(\tau)+k_{\infty}(\nu)\,r_{*}(\tau)=\omega\tau+\nu\left[t(\tau)-r_{*}(\tau)\right]+\Delta(\nu)\,r_{*}(\tau),\quad\Delta(\nu)=\nu-k_{\infty}(\nu).$ (53)
Summing over the three physical polarizations can be organized with the projector $\Pi_{\mu\nu}=-\eta_{\mu\nu}+k_{\mu}k_{\nu}/m_{V}^{2}$ (cf. (33)),
$\sum_{\lambda=1}^{3}\big|\hat{d}^{\mu}\varepsilon^{(\lambda)}_{\mu}\big|^{2}=\hat{d}^{\mu}\hat{d}^{\nu}\,\Pi_{\mu\nu}(\nu,m_{V}),$ (54)
@@
$P_{\text{exc}}^{\text{(dip)}}(\nu,\ell)=\sum_{m,\lambda}\Big|\mathcal{A}^{\text{(dip)}}_{\nu\ell m\lambda}\Big|^{2}=g_{d}^{2}\,|d_{eg}|^{2}\,\Bigg|\int d\tau\,\frac{\Omega(\tau,\nu)\,e^{\,i\Phi(\tau,\nu)}}{\sqrt{k_{\infty}(\nu)}\,r(\tau)}\,\mathcal{D}_{\ell m}(\tau,\nu)\Bigg|^{2},$ (55)
@@
$\mathcal{D}_{\ell m}(\tau,\nu)\equiv Y_{\ell m}^{\ast}\!\left(\theta(\tau),\phi(\tau)\right)\,\left[\hat{d}^{\mu}\hat{d}^{\nu}\,\Pi_{\mu\nu}(\nu,m_{V})\right]^{1/2}.$ (56)
Relative to the current coupling in (44)-(45), the dipole response carries an extra factor $\Omega(\tau,\nu)$ arising from the electric field $E_{\mu}=F_{\mu\nu}u^{\nu}$. In the large-$\omega$ regime, the stationary-phase point lies on the near-horizon segment where the rapidly varying phase is controlled by $\nu[t(\tau)-r_{*}(\tau)]$, guaranteeing the same Planckian detailed-balance factor as in the scalar derivation, while the $\Omega(\tau,\nu)$ and projector in (56) modify only the smooth prefactor and the threshold behavior inherited from $k_{\infty}(\nu)$.
IV.3 Counter-rotating channel and selection rules
With the Boulware vacuum as the in-state and atoms initially in $|g\rangle$, only the counter-rotating processes (atom excitation accompanied by field emission) survive at first order. The rotating processes either (i) excite the atom while annihilating a quantum $(a_{\nu\ell m\lambda})$, which vanishes on $|0_{B}\rangle$, or (ii) de-excite the atom while creating a quantum-irrelevant for atoms prepared in $|g\rangle$. This is the same mechanism emphasized in Scully et al. Scully:2017utk for the scalar case and is the operative origin of emission-before-absorption.
For the current coupling (35) the relevant vertex on $|g;0_{B}\rangle$ is
$\mathcal{V}^{\text{(mono)}}_{\rm cr}(\tau)\;=\;g_{V}\,m_{eg}\,e^{+i\omega\tau}\,u^{\mu}(\tau)\,A_{\mu}^{(+)\,\dagger}\!\left(x(\tau)\right),$ (57)
and for the dipole coupling (46)
$\mathcal{V}^{\text{(dip)}}_{\rm cr}(\tau)\;=\;g_{d}\,d_{eg}\,e^{+i\omega\tau}\,\hat{d}^{\mu}(\tau)\,F_{\mu\nu}^{(+)\,\dagger}\!\left(x(\tau)\right)\,u^{\nu}(\tau),$ (58)
where the $(+)$ parts denote the positive-Killing-frequency mode operators. The corresponding amplitudes have already been reduced to the single worldline integrals in (40) and (52); their phases are governed by $\nu[t(\tau)-r_{*}(\tau)]$, which is what ultimately produces the Planckian detailed-balance factor as in the scalar derivation.

### v2

@@ IV Detector models and atom-field interaction
We model the atom as a two-level system with gap $\omega>0$ moving on the infalling geodesic specified in Sec. II. The Proca field is expanded in properly normalized outgoing Schwarzschild modes as in Sec. III. Interaction is treated perturbatively in the coupling constants. We consider two couplings: a charged-monopole (current) coupling, where the atomic monopole couples to $u^{\mu}A_{\mu}$ along the trajectory, and a physical dipole coupling to the local electric field. In each case, excitation of a ground-state atom in the Boulware vacuum is mediated by the counterrotating channel (atom excitation + field emission). The resulting amplitudes reduce to a single worldline integral with the near-horizon phase $\nu[t(\tau)-r_{*}(\tau)]$.
@@ IV.1 Charged-monopole (current) coupling
$L_{\text{int}}(\tau)=g_{V}\,m(\tau)\,u^{\mu}(\tau)A_{\mu}\!\left(x(\tau)\right),$ (IV.1)
@@
$m(\tau)=m_{eg}\,e^{+i\omega\tau}\,\sigma_{eg}+m_{ge}\,e^{-i\omega\tau}\,\sigma_{ge},\quad m_{ge}=m_{eg}^{\ast},$ (IV.2)
@@
$A_{\mu}(x)$ $=\sum_{\ell m}\sum_{\lambda=1}^{3}\int_{m_{V}}^{\infty}\!d\nu\;\Big[a_{\nu\ell m\lambda}\,\mathcal{A}^{(\lambda)}_{\mu;\nu\ell m}(x)$
$+a_{\nu\ell m\lambda}^{\dagger}\,\mathcal{A}^{(\lambda)\,*}_{\mu;\nu\ell m}(x)\Big],$ (IV.3)
@@
$\mathcal{A}^{(\lambda)}_{\mu;\nu\ell m}(x)$ $\sim\frac{1}{r}\,\frac{\varepsilon^{(\lambda)}_{\mu}(\nu\,\hat{\mathbf{k}})}{\sqrt{k_{\infty}(\nu)}}\,Y_{\ell m}(\theta,\phi)\,e^{-i\nu t}\,e^{-ik_{\infty}(\nu)r_{*}},$
$k_{\infty}(\nu)$ $=\sqrt{\nu^{2}-m_{V}^{2}}.$ (IV.4)
Near the horizon each outgoing mode reduces to the universal WKB form $e^{-i\nu[t-r_{*}]}$, consistent with (II.15) and the sign convention fixed earlier.
Starting from the ground state $|g;0_{B}\rangle$, the only first-order process that excites the atom in the Boulware vacuum is the counterrotating one, $|g;0_{B}\rangle\to|e;1_{\nu\ell m\lambda}\rangle$. The corresponding amplitude for a selected cavity mode $(\nu,\ell,m,\lambda)$ is
$\mathcal{A}^{\text{(mono)}}_{\nu\ell m\lambda}=-\,ig_{V}m_{eg}\!\int_{-\infty}^{+\infty}\!d\tau\;e^{+i\omega\tau}\,u^{\mu}(\tau)\,\mathcal{A}^{(\lambda)\,*}_{\mu;\nu\ell m}\!\left(x(\tau)\right).$ (IV.5)
Inserting (IV.4) and restricting to the outgoing branch selected by the cavity, we obtain a one-dimensional integral of the form
$\mathcal{A}^{\text{(mono)}}_{\nu\ell m\lambda}$ $=-\,ig_{V}m_{eg}\!\int d\tau\;\frac{u^{\mu}(\tau)\,\varepsilon^{(\lambda)}_{\mu}(\nu\,\hat{\mathbf{k}})}{\sqrt{k_{\infty}(\nu)}\,r(\tau)}\,$
$\times Y_{\ell m}^{\ast}\!\left(\theta(\tau),\phi(\tau)\right)\,e^{\,i\Phi(\tau,\nu)},$ (IV.6)
@@
$\Phi(\tau,\nu)\equiv\omega\tau+\nu\,t(\tau)+k_{\infty}(\nu)\,r_{*}(\tau).$ (IV.7)
Using the geodesic relations (II.5)–(II.8) and the near-horizon WKB replacement $k_{\infty}(\nu)\to\nu$ inside the rapidly varying part, we may equivalently write
$\Phi(\tau,\nu)$ $=\omega\tau+\nu\,\left[\,t(\tau)-r_{*}(\tau)\,\right]+\Delta(\nu)\,r_{*}(\tau),$
$\Delta(\nu)$ $\equiv\nu-k_{\infty}(\nu),$ (IV.8)
where $\Delta(\nu)$ is smooth and slowly varying. The polarization contraction can be organized with the projector defined in (III.17). Summing over the three Proca polarizations gives the smooth weight
$\sum_{\lambda=1}^{3}\big|u^{\mu}\varepsilon^{(\lambda)}_{\mu}\big|^{2}$ $=u^{\mu}u^{\nu}\left(-\eta_{\mu\nu}+\frac{k_{\mu}k_{\nu}}{m_{V}^{2}}\right)$
$=-\,u^{\mu}u_{\mu}+\frac{(u\!\cdot\!k)^{2}}{m_{V}^{2}}=-1+\frac{\Omega^{2}(\tau,\nu)}{m_{V}^{2}},$ (IV.9)
where $\Omega(\tau,\nu)=-k_{\mu}u^{\mu}$ is the comoving frequency along the worldline, equal to the redshift-Doppler factor given in (II.15) for the near-horizon outgoing branch.
@@
$P_{\text{exc}}^{\text{(mono)}}(\nu,\ell)$ $=\sum_{m,\lambda}\Big|\mathcal{A}^{\text{(mono)}}_{\nu\ell m\lambda}\Big|^{2}$
$=g_{V}^{2}\,|m_{eg}|^{2}\,\Bigg|\int d\tau\,\frac{e^{\,i\Phi(\tau,\nu)}}{\sqrt{k_{\infty}(\nu)}\,r(\tau)}\,\mathcal{S}_{\ell m}(\tau,\nu)\Bigg|^{2},$ (IV.10)
@@
$\mathcal{S}_{\ell m}(\tau,\nu)\equiv Y_{\ell m}^{\ast}\!\left(\theta(\tau),\phi(\tau)\right)\,\left[\,-1+\frac{\Omega^{2}(\tau,\nu)}{m_{V}^{2}}\,\right]^{1/2}$ (IV.11)
collects the angular dependence and the polarization weight. In the large-$\omega$ regime and with the cavity isolating a single outgoing channel, the integral in (IV.10) will be evaluated by stationary phase around the near-horizon segment where $\Phi^{\prime}(\tau,\nu)=0$, with the rapidly varying part governed by $\nu[t(\tau)-r_{*}(\tau)]$ and the smooth prefactor supplied by $\mathcal{S}_{\ell m}$ and $1/r(\tau)$. The appearance of $\Omega(\tau,\nu)$ in (IV.11) traces directly to the current coupling $u\!\cdot\!a$ and encodes the distinctive mass/polarization dependence of the vector response.
@@ IV.2 Physical dipole coupling
$L_{\text{int}}(\tau)\;=\;-\,d^{\mu}(\tau)\,F_{\mu\nu}\!\left(x(\tau)\right)\,u^{\nu}(\tau),\qquad u_{\mu}d^{\mu}(\tau)=0,$ (IV.12)
@@
$d^{\mu}(\tau)=d_{eg}\,e^{+i\omega\tau}\,\hat{d}^{\mu}\,\sigma_{eg}+d_{ge}\,e^{-i\omega\tau}\,\hat{d}^{\mu}\,\sigma_{ge},$
$\hat{d}^{\mu}\hat{d}_{\mu}=-1,\quad u_{\mu}\hat{d}^{\mu}=0.$ (IV.13)
Using the properly normalized outgoing modes from Sec. III, the field strength associated with a single mode $(\nu,\ell,m,\lambda)$ is
$F_{\mu\nu}^{(\lambda)}(x)\;=\;i\left[k_{\mu}\,\mathcal{A}^{(\lambda)}_{\nu;\nu\ell m}(x)-k_{\nu}\,\mathcal{A}^{(\lambda)}_{\mu;\nu\ell m}(x)\right],$ (IV.14)
where $\mathcal{A}^{(\lambda)}_{\mu;\nu\ell m}$ has the asymptotics in (IV.4) and $k^{\mu}=(\nu,\mathbf{k})$ with $k_{\infty}=\sqrt{\nu^{2}-m_{V}^{2}}$. Contracting with $u^{\nu}$ defines the electric field in the atom’s instantaneous rest frame [43, 21],
$E_{\mu}\;\equiv\;F_{\mu\nu}\,u^{\nu}\;=\;i\left[-(u\!\cdot\!k)\,\mathcal{A}_{\mu}+(u\!\cdot\!\mathcal{A})\,k_{\mu}\right].$ (IV.15)
@@
$E_{\mu}=i\frac{Y_{\ell m}}{r\sqrt{k_{\infty}(\nu)}}\Big[k_{\mu}(\varepsilon\cdot u)-\Omega(\tau,\nu)\varepsilon_{\mu}\Big]e^{-i\nu t}e^{-ik_{\infty}r_{*}}+\cdots,$ (IV.16)
with $\Omega=-k\cdot u$. The counterrotating channel $|g;0_{B}\rangle\!\to\!|e;1_{\nu\ell m\lambda}\rangle$ gives the first-order amplitude
$\mathcal{A}^{\text{(dip)}}_{\nu\ell m\lambda}=-\,ig_{d}d_{eg}\!\int d\tau\;\hat{d}^{\mu}\,E_{\mu}^{(\lambda)\,*}\!\left(x(\tau)\right)\,e^{+i\omega\tau},$ (IV.17)
which, after inserting (IV.16) and using the cavity-selected outgoing branch, reduces to
$\mathcal{A}^{\text{(dip)}}_{\nu\ell m\lambda}$ $=-g_{d}d_{eg}\int d\tau\frac{e^{i\Phi(\tau,\nu)}}{\sqrt{k_{\infty}(\nu)}r(\tau)}$
$\times\left[\Omega(\tau,\nu)\hat{d}\cdot\varepsilon^{(\lambda)}-(\hat{d}\cdot k)(\varepsilon^{(\lambda)}\cdot u)\right]Y_{\ell m}^{\ast}(\theta(\tau),\phi(\tau)),$ (IV.18)
with the same phase used in (IV.7)–(IV.8),
$\Phi(\tau,\nu)$ $=\omega\tau+\nu\,t(\tau)+k_{\infty}(\nu)\,r_{*}(\tau)$
$=\omega\tau+\nu\left[t(\tau)-r_{*}(\tau)\right]+\Delta(\nu)\,r_{*}(\tau),$
$\Delta(\nu)$ $=\nu-k_{\infty}(\nu).$ (IV.19)
Summing over the three physical polarizations can be organized with the projector $\Pi_{\mu\nu}=-\eta_{\mu\nu}+k_{\mu}k_{\nu}/m_{V}^{2}$ (cf. (III.17)),
$\sum_{\lambda=1}^{3}\big|\hat{d}^{\mu}\varepsilon^{(\lambda)}_{\mu}\big|^{2}=\hat{d}^{\mu}\hat{d}^{\nu}\,\Pi_{\mu\nu}(\nu,m_{V}),$ (IV.20)
@@
$P_{\text{exc}}^{\text{(dip)}}(\nu,\ell)$ $=\sum_{m,\lambda}\Big|\mathcal{A}^{\text{(dip)}}_{\nu\ell m\lambda}\Big|^{2}=g_{d}^{2}\,|d_{eg}|^{2}\,$
$\times\Bigg|\int d\tau\,\frac{\Omega(\tau,\nu)\,e^{\,i\Phi(\tau,\nu)}}{\sqrt{k_{\infty}(\nu)}\,r(\tau)}\,\mathcal{D}_{\ell m}(\tau,\nu)\Bigg|^{2},$ (IV.21)
@@
$\mathcal{D}_{\ell m}(\tau,\nu)\equiv Y_{\ell m}^{\ast}\!\left(\theta(\tau),\phi(\tau)\right)\,\left[\hat{d}^{\mu}\hat{d}^{\nu}\,\Pi_{\mu\nu}(\nu,m_{V})\right]^{1/2}.$ (IV.22)
Relative to the current coupling in (IV.10)–(IV.11), the dipole response carries an extra factor $\Omega(\tau,\nu)$ arising from the electric field $E_{\mu}=F_{\mu\nu}u^{\nu}$. In the large-$\omega$ regime, the stationary-phase point lies on the near-horizon segment where the rapidly varying phase is controlled by $\nu[t(\tau)-r_{*}(\tau)]$, guaranteeing the same Planckian detailed-balance factor as in the scalar derivation, while the $\Omega(\tau,\nu)$ and projector in (IV.22) modify only the smooth prefactor and the threshold behavior inherited from $k_{\infty}(\nu)$.
IV.3 Counterrotating channel and selection rules
With the Boulware vacuum as the in-state and atoms initially in $|g\rangle$, only the counterrotating processes (atom excitation accompanied by field emission) survive at first order. The rotating processes either (i) excite the atom while annihilating a quantum $(a_{\nu\ell m\lambda})$, which vanishes on $|0_{B}\rangle$, or (ii) deexcite the atom while creating a quantum-irrelevant for atoms prepared in $|g\rangle$. This is the same mechanism emphasized in Scully et al. [44] for the scalar case and is the operative origin of emission-before-absorption.
For the current coupling (IV.1) the relevant vertex on $|g;0_{B}\rangle$ is
$\mathcal{V}^{\text{(mono)}}_{\rm cr}(\tau)\;=\;g_{V}\,m_{eg}\,e^{+i\omega\tau}\,u^{\mu}(\tau)\,A_{\mu}^{(+)\,\dagger}\!\left(x(\tau)\right),$ (IV.23)
and for the dipole coupling (IV.12)
$\mathcal{V}^{\text{(dip)}}_{\rm cr}(\tau)\;=\;g_{d}\,d_{eg}\,e^{+i\omega\tau}\,\hat{d}^{\mu}(\tau)\,F_{\mu\nu}^{(+)\,\dagger}\!\left(x(\tau)\right)\,u^{\nu}(\tau),$ (IV.24)
where the $(+)$ parts denote the positive-Killing-frequency mode operators. The corresponding amplitudes have already been reduced to the single worldline integrals in (IV.6) and (IV.18); their phases are governed by $\nu[t(\tau)-r_{*}(\tau)]$, which is what ultimately produces the Planckian detailed-balance factor as in the scalar derivation.